OpenStudy (idkanything):

y=-(6+x) any ideas???

OpenStudy (eyust707):

graph?

OpenStudy (idkanything):

yes :p lol

OpenStudy (eyust707):

well y = -6 -x y = -1x -6

OpenStudy (eyust707):

do you know y=mx+b?

OpenStudy (idkanything):

I've heard of it but i dont remember :(

OpenStudy (eyust707):

y=mx+ b is the equation of a line. m and b are both numbers. m equals the slope of the line b equals the y-intercept

OpenStudy (eyust707):

we can use these two numbers to graph the line

OpenStudy (idkanything):

oh ok i get u

OpenStudy (idkanything):

but how can we use those two numbers to graph the line?

OpenStudy (eyust707):

so whats the slope?

OpenStudy (idkanything):

the slope is -1x i think

OpenStudy (eyust707):

remember, mis just the number in front.. so slope is -1

OpenStudy (idkanything):

oh ok

OpenStudy (eyust707):

so slope = -1

OpenStudy (eyust707):

what is the y intercept?

OpenStudy (idkanything):

-6

OpenStudy (eyust707):

yep

OpenStudy (eyust707):

so we have a line with the slope of -1 that crosses y at -6

OpenStudy (eyust707):



OpenStudy (eyust707):

try to draw the line in

OpenStudy (eyust707):

i labeled the -6 intercept for you.. we know it crosses through that point

OpenStudy (idkanything):